OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = 1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (campbell_st):

the slope of your line is m = 1/3 so the slope of the perpendicular is -1/m then when you have that use the point slope form y = mx + b and substitute x = 4 and y = -2 to find b

OpenStudy (anonymous):

confused :/

OpenStudy (campbell_st):

ok... so if the slope of you given line is m = 1/3 what is -1/m

OpenStudy (anonymous):

1/-3?

OpenStudy (campbell_st):

not quite... you need the negative reciprocal of 1/3 ...

OpenStudy (anonymous):

-3/1?

OpenStudy (campbell_st):

correct or make it simple the slope of the perpendicular line is -3 so the line looks like y = -3x + b to find b, use the point (4, -2) and substitute x = 4 and y = -2 to get the value of b

OpenStudy (anonymous):

ugh what do I do from there?

OpenStudy (campbell_st):

ok... replace x with 4 and y with -2 so you have \[-2 = -3 \times 4 + b\] you need to calculate the value of b

OpenStudy (anonymous):

how do you calculate the value of B?

OpenStudy (anonymous):

never mind I got it

OpenStudy (campbell_st):

ok... you if you simplify the equation above you ge \[-2 = -12 + b\] whats the value of b..?

OpenStudy (anonymous):

10

OpenStudy (campbell_st):

so the equation is y = -3x + 10

OpenStudy (anonymous):

yeess :) thanks!